The probability that a randomly selected 1-year-old male stink bug will live to be 2 years old is 0.95028 (a) What is the probability that two randomly selected 1-year-old male stink bugs will live to be 2 years old? (b) What is the probability that eight randomly selected 1-year-old male stink bugs will live to be 2 years old? (c) What is the probability that at least one of eight randomly selected 1-year-old male stink bugs will not live to be 2 years old? Would it be unusual if at least one of eight randomly selected 1-year-old male stink bugs did not live to be 2 years old? a) The probability that two randomly selected 1-year-old male stink bugs will live to be 2 years old is Round to five decimal places as needed.) (b) The probability that eight randomly selected 1-year-old male stink bugs will live to be 2 years old is Round to five decimal places as needed.) (c) The probability that at least one of eight randomly selected 1-year-old male stink bugs will not live to be 2 years old is (Round to five decimal places as needed.) Would it be unusual if at least one of eight randomly selected 1-year-old male stink bugs did not live to be 2 years old? because the probability of this happening is 0.05
Answer #1

a)
probabiilty = 0.95028 * 0.95028 = 0.90303


b)
probability = 0.95028^8 = 0.66499

c)
probability = 1 - 0.66499 = 0.33501

No, because probability of this happening is greater than 0.05